OpenStudy (anonymous):

Find the dimensions of a rectangle with area 2500 m^2 whose perimeter is as small as possible. List the dimensions in non-decreasing order.

OpenStudy (anonymous):

i got x value to be sqrt(1250) and y value to be 2500/sqrt(1250)

OpenStudy (anonymous):

Is this calc 3? or calc 1?

OpenStudy (anonymous):

\[ A=lw=2500\implies l=2500/w \]So \[ P=2w+2(2500/w)= 2w+5000/w \]We take derivative\[ P'=2+\frac{-5000}{w^2} \]To minimize, it must be 0, so \[ 0=2-5000/w^2\implies -2=-5000/w^2\implies w^2=\frac{-2}{-5000} \]

OpenStudy (anonymous):

So, based on what I'm getting. It looks like your answer is correct.

OpenStudy (anonymous):

By the way \[ \sqrt{2500}=\sqrt{25\cdot 100}=\sqrt{25}\cdot \sqrt{100}=5\cdot 10=50 \]

OpenStudy (anonymous):

So you should have x=50, and y=2500/50=50

OpenStudy (anonymous):

I guess the only way to minimize a rectangle's perimeter is to make it a square.